Operation Summary for January 11, 2010 - January 13, 2010 |
Monday Jan. 11 The day shift began with Operations monitoring store 7494 (110.8E30) and the stash (194.6E10), with the Antiproton Source (Pbar) stacking (17.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7494 at 3 PM. Operators prepared the CDF and D0 collision hall for accesses. The test stand I- source was replaced; the new source was warmed up through the end of the shift. Booster personnel conducted studies. During the evening shift, Operations reported that CDF and D0 came out of access at 5:40 PM. Operations put the TeV through a dry squeeze. A Linac RF station (LRF2) tripped off and would not reset. Linac personnel came in to work on the problem, which lasted almost until the end of the shift. Operations made adjustments to the H- Source as needed. Tuesday Jan. 12 The midnight shift began with Operations monitoring the stash (455.3E10), with Pbar stacking (17.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations started shift by checking the TeV orbits. Operations put the TeV into HEP shot setup at 1:33 AM and established store 7499 at 2:52 AM with an average initial luminosity (AIL) of 263.6E30. At the end of the shift, Operations reported that the output of the Booster Chopper (B:CHOP) was dropping. During the day shift, Booster personnel traced the chopper fault to the "on" tube high voltage cable; they reterminated the cable. Operations resumed sending Booster beam around 9:45 AM. Booster personnel conducted studies during this shift. During the evening shift, around 4 PM, the TeV magnet tilt meters recorded the 7.0 magnitude earthquake that hit Haiti. Operations terminated store 7499 at 5:54 PM, and then put the TeV into shot setup. Operations established store 7500 at 8:11 PM with an AIL of 280.67E30. At 11:11 PM, LRF2 went down. Operators were unable to recover the station. They called in Linac technicians to make repairs. The repairs lasted through to the end of the shift.
Wednesday Jan. 13 The midnight shift began with Operations monitoring store 7500 (141.31E30), the stash (130.2E10), and the Pbar stack (18.7E10). NuMI and MiniBooNE were off due to LRF2 problems. During the midnight shift, Linac systems people worked on LRF2 all shift; there was no stacking or beam to NUMI and MiniBooNE. The Plans for Wednesday The plans for today are to bring LRF2 back online. The Run Coordinator will allow maintenance work and accesses that don’t interfere with the stash or stack. |
|
More Information Comments and Suggestions |
| last modified 02/03/2010 email Fermilab |
| Security, Privacy, Legal | |